2x(squared) + 42x + 441 = 2601
Answer by ankor@dixie-net.com(22740)

2x^2 + 42x + 441 = 2601
2x^2 + 42x + 441 - 2601 = 0
2x^2 + 42x - 2160 = 0
simplify, divide by 2:
x^2 + 21x - 1080 = 0
You can use the quadratic formula to find x, but this will factor to:
(x+45)(x-24) = 0
Two solutions
x = -45
x = 24